A coin bank contains nickels, dimes, and quarters totaling $5.45. If there are twice as many quarters as dimes and 11 more nickels than quarters, how many of each coin are in the bank?
Answer by stanbon(75887)

A coin bank contains nickels, dimes, and quarters totaling $5.45. If there are twice as many quarters as dimes and 11 more nickels than quarters, how many of each coin are in the bank?
--------------------------------------
Value Equation: 5n + 10d + 25q = 545 cents
Quantity Eq...: q = 2d
Quantity Eq...: n = q + 11
-------------------------------------------
Rearrange the 2nd equation:
d = (1/2)q
-------------------------------------------
Substitute to solve for "q":
5(q+11) + 10(1/2)q + 25q = 545
5q + 55 + 5q + 25q = 545
35q = 490
q = 14 (# of quarters)
-------------
Since n = q + 11, n = 25 (# of nickels)
Since d = (1/2)q, d = 7 (# of dimes)
==================================
Cheers,
stan H.